OpenStudy (anonymous):

Solve 2(x - 4) - 6(2x - 3) = 0 A. x = -4 B. x = -1 C. x = 0 D. x = 1

OpenStudy (callisto):

First expand the terms 2(x - 4) - 6(2x - 3) = 0 2x -2(4) - 6(2x) -6(-3) =0 2x -8 -12x +18 =0 When simplify the terms' 2x -12x -8 +18 =0 x( 2-12) +10 =0 -10x +10 =0 Minus both sides by 10 -10x +10 -10 = 0=10 -10x =-10 Divide both sides by -10 Can you work from here and the the answer?

OpenStudy (anonymous):

I have a recommendation of skills for exams. As some values of x are given, if you really do not know how to find z, you can put all options into the question one by one. If you can get the result which is zero. That's the answer. But, it is better for you to learn the method suggested by callisto

OpenStudy (callisto):

Somehow, for solving the equation like this with single unknown, it's faster to solve it and entering all the answers to your calculator...